OpenStudy (anonymous):

When there is a negativepostive in a (y=mx+-b or y=m-+b) does the negative or positive sign take over?

OpenStudy (anonymous):

The negative sign takes over? The question is confusing.

OpenStudy (anonymous):

equation is y =mx+b

OpenStudy (anonymous):

for ex if you have (-9,-3) and a slope of a -6. you have to put it into a y=m(x-x1)+y1

OpenStudy (anonymous):

so it would be y=-6(x+9)+3 right?

OpenStudy (anonymous):

Wait what, why'd you put x+9?

OpenStudy (anonymous):

and +3?

OpenStudy (anonymous):

because you plug them in from the points

OpenStudy (anonymous):

How'd you know x and b?

OpenStudy (anonymous):

y=-6(x+9)+3 y=-6x-54+3 y=-6x-51

OpenStudy (anonymous):

No i was just wondering if those points were given or not.

OpenStudy (anonymous):

they are give

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

i just wanted to know if the 3 in y=-6(x+9)+3 would be positive or negative. you know what i mean?

OpenStudy (anonymous):

If you're y intercept is negative then it will be negative, if it's positive then it will be positive.

OpenStudy (anonymous):

your*